OpenStudy (anonymous):

If a ≠ b and 1/x + 1/a = 1/b, then x =

OpenStudy (anonymous):

then flip both sides to get \(x\)

OpenStudy (anonymous):

you good from there or no?

OpenStudy (anonymous):

So when you subtract a from b, it just turns it into /ab.? you just put the 2 beside each other.?

OpenStudy (anonymous):

well it is a good thing you are confused, because i see that i made a mistake let me start again \[\frac{1}{x}+\frac{1}{a}=\frac{1}{b}\] lets do it correctly this time you want to get \(x\) by itself, so the first thing to do is subract \(\frac{1}{a}\) from both sides and write \[\frac{1}{x}=\frac{1}{b}-\frac{1}{a}\]

OpenStudy (anonymous):

then you have to actually do the subtraction on the right, i.e write \[\frac{1}{b}-\frac{1}{a}\] as a single fraction

OpenStudy (anonymous):

when you subtract you get \[\frac{1}{x}=\frac{a-b}{ab}\] then flip both sides to get \[x=\frac{ab}{a-b}\]

OpenStudy (anonymous):

"Good thing you are confused" haha, but why would I need to flip it.?

OpenStudy (anonymous):

because the left hand side was \(\frac{1}{x}\) and you wanted to solve for \(x\)

OpenStudy (anonymous):

so how to you get from \(\frac{1}{x}\) to \(x\) ? by flipping it the fancy term is "take the reciprocal of both sides"

OpenStudy (anonymous):

Ahhh.! reciprocal. thanks.!

OpenStudy (anonymous):

yw